Calculate the enthalpy change for the following chemical reaction:



CS2 (l) + 3O2 (g) —> CO2 (g) + 2SO2 (g)



1) C (s) + O2 (g) —> CO2 (g)


ΔH°= -393.5 kJ



2) S (s) + O2 (g) —> SO2 (g)


ΔH°= -296.8 kJ



3) C (s) + 2S (s) —> CS2 (l)


ΔH°= +87.9 kJ

1
Expert's answer
2022-05-06T13:39:04-0400

eq - equation

(1st eq): C(s) + O2(g) → CO2(g), ∆H = −393.5 kJ/mol

(2nd eq): S(s) + O2(g) → SO2(g), ∆H = −296.8 kJ/mol

(3d eq): C(s) + 2S(s) → CS2(l), ∆H = +87.9 kJ/mol

 

CS2(l) + 3O2(g) → CO2(g) + 2SO2(g), ∆Hx = ???

 

Solution:

According to Hess's law, the heat of reaction depends upon initial and final conditions of reactants and does not depend of the intermediate path of the reaction.

 

1) Modify the three given equations to get the target equation:

(1st eq): do nothing. We need one mole of CO2(g) on the product side and that's what we have.

(2nd eq): multiply by two. We need two mole of SO2(g) on the product side.

(3d eq): flip it so as to put CS2(l) on the reactant side.


2) Rewrite all three equations with the changes made (including changes in their enthalpy):

(1st eq): C(s) + O2(g) → CO2(g), ∆H1 = −393.5 kJ/mol

(2*nd eq): 2S(s) + 2O2(g) → 2SO2(g), ∆H2 = 2 × (−296.8) = −593.6 kJ/mol

(3*deq): CS2(l) → C(s) + 2S(s), ∆H3 = (−1) × (+87.9) = −87.9 kJ/mol


3) Cancel out the common species on both sides:

C(s) ⇒ (1st equation) & (3*d equation)

2S(s) ⇒ (2*nd equation) & (3*d equation)


Thus, adding modified equations and canceling out the common species on both sides, we get:

CS2(l) + 3O2(g) → CO2(g) + 2SO2(g)


4) Add the ΔH values of (1st), (2*nd eq) and (3*d) equations to get your answer:

∆Hx = ∆H1 + ∆H2 + ∆H3 = (−393.5 kJ/mol) + (−593.6 kJ/mol) + (−87.9 kJ/mol) = −1075 kJ/mol

∆Hx = −1075 kJ/mol

Therefore, the enthalpy change for the following chemical reaction is −1075 kJ/mol


Answer: The enthalpy change for the reaction is −1075 kJ/mol